OpenStudy (anonymous):

How much does Yusuf need to invest at a rate of 6% in order to earn $150 interest in 5 years?

OpenStudy (anonymous):

I=P*R*T

OpenStudy (anonymous):

Hint: Solve for P P=IRT I = 150 R = .06 T = 5

OpenStudy (mathstudent55):

@Trai7 When you write "solve for P" you should mean to solve the equation correctly for P, not just rearrange it incorrectly with P equaling something.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Im so confused :/

OpenStudy (anonymous):

I keep coming up with 500$. I think that that is correct.

OpenStudy (mathstudent55):

\(I = PRT\) \(P = \dfrac{I}{RT}\)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok ty

OpenStudy (mathstudent55):

You are correct. \(P = \dfrac{I}{RT} = \dfrac{$150}{0.06 \times 5} = $500\)

OpenStudy (anonymous):

I was right WOO